OpenStudy (anonymous):

There are two consecutive even numbers such that six times the smaller number minus five times the greater number is -2. What is the smaller number?

hero (hero):

Hint: x = smaller number x + 2 = larger number 6x - 5(x+2) = -2

hero (hero):

Solve for x

OpenStudy (waleed_imtiaz):

the numbers are 8 and 10, i.e., 8*6-10*5=-2 so.. 8 is the smallest number of the two...... :)

OpenStudy (waleed_imtiaz):

Just start from the lower consecutive even numbers..... try for 2 or 3 times will give U the answer..... Otherwise , the least possible numbers which give =-2 difference are 8 and 10. ANd I don't think any explanation is needed else........ THanks.... Otherwise.. I like ur justification @Hero

hero (hero):

You're not just supposed to give the answer @waleed_imtiaz, that's the whole point. This is Open Study not Open Question and Answer.

OpenStudy (waleed_imtiaz):

THank you very much @Hero for ur concern..... But its not that I wasn't confirmed about my answer. I will give the answer after knowing that my explanation and calculation is correct. Em not just making guess here... Okay... Sorry, If u don't like my explanation about this question. Everyone has his own point of view....

OpenStudy (waleed_imtiaz):

Why I need to jump into the equations always if I have another simple way to do it... :)